OpenStudy (anonymous):

If I buy 2 apples and 6 oranges, I will spend $5.60. However, if I buy 6 apples and 2 oranges, I will spend $4.80. How much does one orange cost? A. $0.55 B. $0.70 C. $0.75 D. $0.60

OpenStudy (cwrw238):

this is another problem solvable by systems of equations 2x + 6y = 5.60 6x + 2y = 4.80 (x=apples y = oranges)